Um.... hello ^_^ I was absent by the day this was taught.....and no one can help me T_T.... will you??? :)
How many quarts of pure acid must be added to 5 quarts of a 30% acid solution to produce a 50% acid solution ?

Answer by mananth(16946)

--- percent ---------------- Amount
acid I 30 ---------------- 5 quarts
Pure acid 100 ---------------- x quarts
Mixture 50 ---------------- 5 + x quarts

30 * 5 + 100 x = 50 ( 5 + x )
150 + 100 x = 250 + 50 x
100 x -50 x = -150 + 250
50 x = 100
/ 50
x = 2 quarts Pure acid

m.ananth@hotmail.ca